OpenStudy (anonymous):

Can someone explain how to find the inverse of a function like F(x)=1/3x+2

OpenStudy (er.mohd.amir):

let y=1/(3x+2) (3x+2)y=1 3xy+2y=1 3xy=1-2y x=(1-2y)/3y

OpenStudy (er.mohd.amir):

that is inverse is=(1-2x)/3x

OpenStudy (anonymous):

How would I show they are inverses using function composition? Thanks for the help!!